# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ Abstract
Oriented object detection is one of the most fundamental and challenging tasks in remote sensing, aiming to locate and classify objects with arbitrary orientations. Recent years have witnessed remarkable progress in oriented object detection using deep learning techniques. Given the rapid development of this field, this paper aims to provide a comprehensive survey of recent advances in oriented object detection. To be specific, we first review the technical evolution from horizontal object detection to oriented object detection and summarize the specific challenges, including feature misalignment, spatial misalignment, and periodicity of angle. Subsequently, we further categorize existing methods into detection framework, oriented bounding box (OBB) regression, and feature representations, and discuss how these methods address the above challenges in detail. In addition, we cover several publicly available datasets and performance evaluation protocols. Furthermore, we provide a comprehensive comparison and analysis of state-of-the-art oriented object detection methods. Toward the end of this paper, we discuss several future directions for oriented object detection.
@@
[…]
With the rapid development of remote sensing (RS) technologies, an increasing number of images with various resolutions and different spectra can be easily obtained by optical satellites or unmanned aerial vehicles (UAVs). Naturally, the research community has an urgent demand to investigate a variety of advanced technologies for automatically and efficiently processing and analyzing massive RS images. As a crucial cornerstone of automatic analysis for RS images, object detection aims to recognize objects of predefined categories from given images and to regress a precise localization of each object instance (Liu et al., 2020; Zou et al., 2023). Nowadays, object detection has served as an essential step for a broad range of RS applications, including intelligent monitoring (Zhao et al., 2018; Salvoldi et al., 2022), precision agriculture (Osco et al., 2021), urban planning (Burochin et al., 2014), port management (Zhang et al., 2021a), and military reconnaissance (Liu et al., 2022c).
Objects in RS images typically appear in arbitrary orientations due to the bird-eye view (BEV), making the general (horizontal) object detection methods unsuitable. Different from general object detection that represents object localization via a horizontal bounding box (HBB), oriented object detection (also called rotated object detection) uses an oriented bounding box (OBB) to tightly pack the oriented object, as shown in Figure 1. The OBB can not only provide orientation information but also locate the object precisely. Consequently, oriented object detection has attracted significant attention, especially in the last five years. While enormous methods exist, a comprehensive survey specifically focused on oriented object detection is still lacking. Given the continued maturity and increasing concerns about this field, this paper attempts to present a thorough analysis of recent efforts and systematically summarize their achievements.
@@ 1.1 Comparisons with Related Surveys
Survey title |
Venue |
Descriptions |
General object detection | ||
A Survey of Deep Learning-Based Object Detection (Jiao et al., 2019) |
ACCESS |
A survey on deep learning frameworks and methods for generic object detection |
Deep Learning for Generic Object Detection: A Survey (Liu et al., 2020) |
IJCV |
@@
A comprehensive survey of the recent progress in generic object detection brought about by deep learning |
Recent advances in deep learning for object detection (Wu et al., 2020) |
Neucom |
@@
A survey focuses on deep learning in generic object detection from detection components, learning strategies, and applications |
Object Detection With Deep Learning: A Review (Zhao et al., 2019) |
TNNLS |
@@
A review on deep learning for generic object detection and other specific subtasks |
Object Detection in 20 Years: A Survey (Zou et al., 2023) |
PROC |
@@
A survey focuses on object detection spanning over 20 years of history |
Specific category | ||
A survey on face detection in the wild: Past, present and future (Zafeiriou et al., 2015) |
CVIU |
A survey on real-world face detection techniques since 2000 |
The Elements of End-to-end Deep Face Recognition: A Survey of Recent Advances (Du et al., 2022) |
ACM |
A survey about each element of end-to-end deep face detection, face alignment, and face representation |
Text Detection and Recognition in Imagery: A Survey (Ye and Doermann, 2015) |
TPAMI |
@@
A survey about methods, sub-problems, and special issues of text detection and recognition |
Text Detection, Tracking, and Recognition in Video: A Comprehensive Survey (Yin et al., 2016) |
TIP |
A survey of text detection, tracking, and recognition methods and systems in the video |
Computer vision and deep learning techniques for pedestrian detection and tracking: A survey (Brunetti et al., 2018) |
Neucom |
A survey on deep learning for pedestrian detection and tracking |
From Handcrafted to Deep Features for Pedestrian Detection: A Survey (Cao et al., 2022) |
TPAMI |
@@
A survey on recent deep features based methods in pedestrian detection |
Specific conditions | ||
Recent advances in small object detection based on deep learning: A review (Tong et al., 2020) |
IVC |
A survey on small object detection using deep learning |
A survey and performance evaluation of deep learning methods for small object detection (Liu et al., 2021a) |
ESWA |
A survey of challenges, solutions, and deep learning technologies for small object detection |
Towards Large-Scale Small Object Detection: Survey and Benchmarks (Cheng et al., 2023) |
TPAMI |
@@
A survey of small object detection and two large-scale small object detection benchmarks under driving scenario and aerial scene |
Few-Shot Object Detection: A Survey (Pannone, 2022) |
ACM |
@@
A survey on few-shot object detection through data augmentation, transfer learning, distance metric learning, and meta-learning |
A Survey of Self-Supervised and Few-Shot Object Detection (Huang et al., 2023) |
TPAMI |
@@
Categorization, review, and comparison for few-shot and self-supervised object detection methods |
Weakly Supervised Object Localization and Detection: A Survey (Zhang et al., 2022a) |
TPAMI |
A survey on weakly supervised object detection, including classic models, off-the-shelf models, and deep weakly supervised approaches |
RS object detection | ||
@@
A survey on object detection in optical remote sensing images (Cheng and Han, 2016) |
ISPRS |
A review on traditional object detection methods in RS images |
Object detection in optical remote sensing images: A survey and a new benchmark (Li et al., 2020) |
ISPRS |
@@
A review on deep learning based horizontal object detection in RS, and a large-scale, publicly available benchmark for RS object detection |
Remote Sensing Object Detection Meets Deep Learning: A Meta-review of Challenges and Advances (Zhang et al., 2023) |
arxiv |
@@
A survey on challenges and advances in RS object detection, including multi-scale object detection, rotated object detection, weak object detection, tiny object detection, and object detection with limited supervision |
@@
Ship detection and classification from optical remote sensing images: A survey (Li et al., 2021) |
CJA |
A survey of RS ship detection schemes from 1978 to 2020 |
@@
Methods for Small, Weak Object Detection in Optical High-Resolution Remote Sensing Images: A survey of advances and challenges (Han et al., 2021b) |
MGRS |
A survey of challenges and recent advances for RS small, weak object detection |
Deep Learning for Unmanned Aerial Vehicle-Based Object Detection and Tracking: A survey (Wu et al., 2022) |
MGRS |
@@
A survey on deep learning approaches in UAV object detection and tracking from static object detection, video object detection, and multiple object detection |
In object detection, quite a number of prominent surveys have been published in recent years, as summarized in Table 1. Many notable surveys focus on generic horizontal object detection that aims to detect horizontal objects in natural scenarios (Jiao et al., 2019; Liu et al., 2020; Wu et al., 2020; Zhao et al., 2019; Zou et al., 2023) and cover various aspects, including deep learning based detection frameworks, training strategies, feature representation, evaluation metrics, and typical applications. There also several efforts focused on a specific category, such as face detection (Zafeiriou et al., 2015; Du et al., 2022), text detection (Ye and Doermann, 2015; Yin et al., 2016), and pedestrian detection (Brunetti et al., 2018; Cao et al., 2022). What’s more, there are also surveys paying their attention to object detection under specific conditions, including small object detection (Tong et al., 2020; Liu et al., 2021a; Cheng et al., 2023), few-shot object detection (Pannone, 2022; Huang et al., 2023), and weakly-supervised object detection (Zhang et al., 2022a). Although a few surveys analyze and summarize RS object detection, they either lack of in-depth analysis for oriented object detection (Cheng and Han, 2016; Li et al., 2020, 2021; Han et al., 2021b; Wu et al., 2022) or categorize the sub-categories belonging to RS object detection as oriented object detection (Zhang et al., 2023). Hence, none is rigorously concentrated on an oriented object detection survey. Given the extremely rapid rate of achievements in this field, a comprehensive survey of previous efforts is crucial for further development. By strictly casting our sight on oriented objects, we thoroughly review hundreds of top journal and conference papers related to oriented object detection in RS images. As a result, a systematic survey of oriented object detection is proposed, which is radically different from previous works. We believe that this survey shall be valuable for researchers to gain a clear understanding of oriented object detection with deep learning.
@@ 1.2 Contributions
The major contributions of this work are summarized as follows:
(1) We briefly review the technical evolution from horizontal object detection to oriented object detection and summarize the main challenges, including feature misalignment, spatial misalignment, and PoA.
(2) We systematically summarize recent advancements in oriented object detection and categorize them in three folds according to their motivation: detection frameworks, OBB regression, and feature representations. Meanwhile, we review dozens of RS datasets related to oriented object detection.
(3) We provide a unified comparison of state-of-the-art methods on several typical datasets and an in-depth analysis of the pros and cons of these methods. Furthermore, we identify several promising directions for future research.
The structure of this paper is organized as follows. We first introduce the development from horizontal object detection to oriented object detection and highlight the major challenges in Section 2. Then, we review deep neural networks (DNN) based detection frameworks in Section 3. Furthermore, we discuss the OBB regression and feature representation in Section 4 and Section 5, respectively. In addition, we summarize other common issues encountered in Rs scenarios in Section 6. After an overview of commonly used datasets is provided in Section 7, we analyze and compare the state-of-the-art methods in Section 8. Finally, we conclude our work and discuss the future directions of oriented object detection in Section 9.
@@ 2 From Horizontal Object Detection to Oriented Object Detection
The early object detection methods rely on handcrafted descriptors (Lowe, 2004; Dalal and Triggs, 2005; Fei-Fei and Perona, 2005; Wright et al., 2009) and machine learning approaches (Cortes and Vapnik, 1995; Blaschke, 2010; Leitloff et al., 2010; Blaschke et al., 2014). These methods often show limited performance due to the weak feature representations. Although lag far behind in accuracy, their instructive insights still have a profound impact on modern detectors, e.g., sliding windows (Viola and Jones, 2001, 2004), hard negative mining, and bounding box regression (Felzenszwalb et al., 2008, 2010). The readers interested in early object detection methods are referred to the recent survey (Cheng and Han, 2016) that provides an in-depth analysis of many classical object detection methods in RS.
The world has witnessed impressive progress in computer vision with the advance of deep neural networks (DNN) since 2012 (Hinton and Salakhutdinov, 2006; LeCun et al., 2015; Chen et al., 2018; He et al., 2016; Krizhevsky et al., 2012, 2017). Benefiting from the continuous improvement of computing resources, DNN can learn high-level patterns from large-scale datasets in an end-to-end fashion. The pioneering studies bring a little glimmer to the object detection field, especially in light of the fact that the performance of handcrafted features-based detectors reached a plateau after 2010. Since then, an increasing number of DNN-based detectors have been proposed and have dominated the state-of-the-art due to their powerful feature representation. Early research in the deep learning era was primarily concerned with designing horizontal object detectors (Girshick et al., 2014; Girshick, 2015; Ren et al., 2017; Liu et al., 2016a; Lin et al., 2020; Redmon et al., 2016; Redmon and Farhadi, 2017; Hei and Jia, 2020; Duan et al., 2019; Zhou et al., 2019; Yang et al., 2019b) for natural scene images taken from a horizontal perspective. In contrast, RS images are typically captured from the BEV, leading to objects appearing in arbitrary orientations. Hence, horizontal object detectors may suffer from the following problems in RS images: (1) HBBs are inclined to contain background, causing DNN to hardly capture discriminative features vital to subsequent regression and classification tasks. (2) The intersection-over-union (IoU) between an HBB and the adjacent HBBs can be very large in dense arrangement scenarios, especially for objects with extremely large aspect ratios, as illustrated in Figure 1(a). Thus, the non-maximum suppression (NMS) technique tends to cause missed detection. Given the above predicament of HBB, OBB is considered more appropriate for RS object detection (Figure 1(b)). To advance object detection research in RS images, a vast collection of large-scale datasets well-annotated with OBBs is released, e.g., DOTA (Xia et al., 2018; Ding et al., 2022), which can serve as an impartial platform for comparing various methods. Furthermore, the challenging datasets can advance object detection toward increasingly complex and tough scenarios.
@@
With the remarkable development of detection frameworks (Ren et al., 2017; Lin et al., 2020; Redmon et al., 2016; Carion et al., 2020), backbone networks (He et al., 2016; Liu et al., 2021b), and robust feature representation (Liu et al., 2022d; Dosovitskiy et al., 2021), the field of object detection has achieved dramatic breakthroughs. Naturally, an intuitive strategy for designing oriented object detectors is to modify representative horizontal object detectors by incorporating an additional channel that predicts the angle (Zhou et al., 2022; Yu and Da, 2023). However, such a straightforward strategy is plagued by several additional challenges, mainly including feature misalignment, spatial misalignment, and periodicity of angle (PoA).
@@
(1) Feature Misalignment. The prevailing generic object detectors typically contain a feature extraction network followed by a detection head, where the latter leverages the feature representations generated from the former to make decisions. However, the feature representations are generally extracted via axis-aligned convolutions, thereby exposing non-negligible misalignment with oriented objects. Such a misaligned feature representation is harmful to oriented objects due to the lack of rotational information, making the detector struggle to identify objects and regress precise OBBs.
@@
(2) Spatial Misalignment. In addition to feature misalignment, the widely used anchor-based detection methods also struggle with spatial misalignment. Generic anchor-based detectors typically use horizontal anchors as priors thereby having limited overlaps to oriented objects, especially for objects with extremely large aspect ratios. This poses a significant challenge to generic label assignment strategies (Ren et al., 2017), which assign positive or negative samples depending on the overlaps. Thus, the naïve anchor generation mechanism commonly fails to provide sufficient positive samples during the whole training.
@@
(3) Periodicity of angle. Current detectors commonly utilize the regression paradigm to locate objects, which has been proven to be effective and yields dramatic achievement. However, the angle regression will suffer from the PoA, causing the angular boundary discontinuity (Yang et al., 2021c, d, 2022; Qian et al., 2021, 2022). Concretely, a small angle difference may cause a large loss change when the angular value approaches the angular boundary range. Thus, PoA can seriously confuse the network, leading to training instability. For more details please refer to Section A of the Appendix.
@@
To cope with the above dilemmas, various works have been made and achieved notable advancements, which mainly follow two paths. One involves constructing well-designed detection frameworks by devising rotated proposal generation networks or refined heads to remedy feature misalignment and spatial misalignment, and the other line of efforts intends to design effective OBB regression through developing new loss functions or OBB representation to eliminate PoA. In addition, high-quality feature representations are crucial for object detection, hence a great deal of effort is concentrated on network designs for better feature representations. We also cover the solutions for several common issues in RS scenarios. In the following, we will systematically analyze and discuss prominent works in oriented object detection from four aspects: Detection Framework (Section 3), OBB Regression (Section 4), Feature Representation (Section 5), and Common Issues (Section 6). Figure 2 shows the taxonomy of oriented object detection methods in this survey.

### v5

@@ Abstract
Oriented object detection is one of the most fundamental and challenging tasks in remote sensing, aiming to locate and classify objects with arbitrary orientations. Recent advancements in deep learning have significantly enhanced the capabilities of oriented object detection. Given the rapid development of this field, this paper presents a comprehensive survey of recent advances in oriented object detection. To be specific, we begin by tracing the technical evolution from horizontal object detection to oriented object detection and highlighting the specific challenges, including feature misalignment, spatial misalignment, and oriented bounding box (OBB) regression problems. Subsequently, we further categorize existing methods into detection framework, OBB regression, and feature representations, and providing an in-depth discussion on how these approaches address the above challenges. In addition, we cover several publicly available datasets and evaluation protocols. Furthermore, we provide a comprehensive comparison and analysis of state-of-the-art methods. Toward the end of this paper, we identify several future directions for oriented object detection.
@@
Oriented object detection, Remote sensing, Deep learning1 Introduction
With the rapid advancement of remote sensing (RS) technologies, an increasing number of images with various resolutions and distinct spectra can be easily obtained by optical satellites or unmanned aerial vehicles (UAVs). Naturally, there is an imperative need within the research community to investigate a variety of advanced technologies for automatically and efficiently processing and analyzing massive RS images. As a pivotal foundation of automatic analysis for RS images, object detection aims to identify objects of predefined categories from given images and to regress a precise localization of each object instance [96, 205]. Currently, object detection constitutes a vital component in a broad range of RS applications, encompassing intelligent monitoring [222], precision agriculture [118], urban planning [4], port management [231], and military reconnaissance [103].
Objects in RS images typically exhibit arbitrary orientations due to the bird-eye view (BEV), making the general (horizontal) object detection methods inadequate. In contrast to general object detection that represents object localization via a horizontal bounding box (HBB), oriented object detection (also called rotated object detection) employs an oriented bounding box (OBB) to tightly pack the oriented object, as shown in Fig. 2. The OBB can not only provide orientation information but also locate the object precisely. Consequently, oriented object detection has attracted considerable attention, especially within the past five years. Although enormous methods exist, a comprehensive survey specifically focused on oriented object detection is still lacking. Given the continued maturity and increasing concerns about this field, this paper seeks to present a thorough analysis of recent efforts and systematically summarize their achievements.
@@ 1.1 Comparisons with Related Surveys
Survey Title | Publication | Descriptions |
@@
Deep Learning for Generic Object Detection: A Survey [96] | IJCV 2020 | A comprehensive survey of the recent progress in generic object detection brought about by deep learning |
@@
Recent Advances in Deep Learning for Object Detection [160] | Neucom 2020 | A survey focuses on deep learning in generic object detection from detection components, learning strategies, and applications |
@@
Object Detection With Deep Learning: A Review [235] | TNNLS 2020 | A review on deep learning for generic object detection and other specific subtasks |
@@
Object Detection in 20 Years: A Survey [205] | PROC 2023 | A survey focuses on object detection spanning over 20 years of history |
@@
Text Detection and Recognition in Imagery: A Survey [184] | TPAMI 2015 | A survey about methods, sub-problems, and special issues of text detection and recognition |
@@
From Handcrafted to Deep Features for Pedestrian Detection: A Survey [15] | TPAMI 2022 | A survey on recent deep features based methods in pedestrian detection |
@@
Towards Large-Scale Small Object Detection: Survey and Benchmarks [21] | TPAMI 2023 | A survey of small object detection and two large-scale small object detection benchmarks under driving scenario and aerial scene |
@@
Few-Shot Object Detection: A Survey [119] | ACM 2022 | A survey on few-shot object detection through data augmentation, transfer learning, distance metric learning, and meta-learning |
@@
A Survey of Self-Supervised and Few-Shot Object Detection [53] | TPAMI 2023 | Categorization, review, and comparison for few-shot and self-supervised object detection methods |
@@
A Survey on Object Detection in Optical Remote Sensing Images [7] | ISPRS 2016 | A review on traditional object detection methods in RS images |
@@
Object Detection in Optical Remote Sensing Images: A Survey and A New Benchmark [99] | ISPRS 2020 | A review on deep learning based horizontal object detection in RS, and a large-scale, publicly available benchmark for RS object detection |
@@
Remote Sensing Object Detection Meets Deep Learning: A Meta-review of Challenges and Advances [234] | GRSM 2023 | A survey on challenges and advances in RS object detection, including multi-scale object detection, rotated object detection, weak object detection, tiny object detection, and object detection with limited supervision |
@@
Ship Detection and Classification from Optical Remote Sensing Images: A Survey [101] | CJA 2021 | A survey of RS ship detection schemes from 1978 to 2020 |
@@
Methods for Small, Weak Object Detection in Optical High-Resolution Remote Sensing Images: A Survey of Advances and Challenges [45] | GRSM 2021 | A survey of challenges and recent advances for RS small, weak object detection |
@@
Deep Learning for Unmanned Aerial Vehicle-Based Object Detection and Tracking: A survey [159] | GRSM 2022 | A survey on deep learning approaches in UAV object detection and tracking from static object detection, video object detection, and multiple object detection |
A Comprehensive Survey of Oriented Object Detection in Remote Sensing Images [153] | ESWA 2023 | A survey on oriented object detection, including rotation invariance, anchor-free mechanism, and loss function. |
In object detection, quite a number of prominent surveys have been published in recent years, as summarized in Tab. 1. Numerous notable surveys concentrate on generic (horizontal) object detection that aims to detect horizontal objects in natural scenarios [96, 160, 235, 205]. These surveys cover various aspects, including deep learning based detection frameworks, training strategies, feature representation, evaluation metrics, and typical applications.
Furthermore, several efforts are devoted to a specific category, such as text detection [184], and pedestrian detection [15]. Additionally, there are also surveys paying their attention to object detection under specific conditions, including small object detection [21], few-shot object detection [119], and weakly-supervised object detection [211].
Although a few surveys analyze and summarize RS object detection, they frequently lack in-depth analysis for oriented object detection [7, 99, 101, 45, 159]. [234] classify the sub-categories belonging to RS object detection as oriented object detection, merely providing a brief introduction to OBB representation and rotation-insensitive feature learning. [153] only focuses on describing the details of previous oriented object detection methods.
Different from previous object detection surveys, which focus on general object detection methods [96, 160, 235, 205], other related fields [184, 15, 21, 119, 211], RS horizontal object detection [7, 99, 101, 45, 159], or a limited number of oriented object detection models [234, 153], this work systematically and comprehensively reviews recent advances in the field. Especially, in comparison to existing surveys related to oriented object detection [234, 153], our survey provide a deeper, more comprehensive dive into this field, a better taxonomy of the literature, and present discussions regarding challenges, comparison, and future directions. It involves in-depth analyses on various aspects, many of which, to the best of our knowledge, have never been discussed in oriented object detection. In particular, we review the technical evolution from horizontal to oriented object detection and summarize the main challenges. We systematically summarize and discuss recent advancements under proposed taxonomies (including detection frameworks, OBB regression, feature representation, and common issues). We provide a comprehensive comparison of state-of-the-art methods on typical datasets, along with an in-depth analysis of the pros and cons of these methods.
1.2 Scope
Fig. 1 shows the increasing number of publications related to ”oriented object detection” or ”rotated object detection” over the past decade or so. Particularly in the last five years, there has been an explosive growth in the number of papers on deep learning based oriented object detection, rendering it impractical to review all of them. Consequently, it is necessary to establish selection criteria to limit our focus to influential papers published in top journals and conferences. Owing to these constraints, we extend our sincere apologies to authors whose works are not included in this paper. It is worth highlighting that we restrict our attention to oriented object detection in single images. Nevertheless, for completeness and better readability, some well-known works on horizontal object detection are also included.
@@ 1.3 Contributions
Our contribution is manifested in four aspects:
(1) A comprehensive review of the technical evolution from horizontal object detection to oriented object detection. Based on the characteristics of RS images and current object detection models, we categorize the main challenges in oriented object detection into four main parts, including feature misalignment, spatial misalignment, OBB regression problems, and common issues.
(2) A thorough taxonomy of oriented object detection methods. Aiming to help researchers gain a deeper understanding of the key features of oriented object detection methods, we categorize and summarize existing oriented object detection methods according to detection frameworks, OBB regression, feature representation, and common issues. .
(3) A comprehensive comparison of state-of-the-art methods. We provide a comprehensive comparison of state-of-the-art methods on typical datasets, along with an in-depth analysis of the pros and cons of these methods. This analysis aims to offer valuable insights into the efficacy and applicability of these methods in addressing the main challenges of oriented object detection.
(4) Overview of open issues and future directions. We thoroughly look over several essential issues, shedding light on potential directions for future research, i.e., lightweight methods, scenario-specific datasets, multi-modal datasets, and large-scale datasets, as well as multi-modal large models.
The structure of this paper is organized as follows. We first introduce the development from horizontal object detection to oriented object detection and highlight the major challenges in Sec. 2. Then, we review deep neural networks (DNN) based detection frameworks in Sec. 3. Furthermore, we discuss the OBB regression and feature representation in Sec. 4 and Sec. 5, respectively. In addition, we summarize other common issues encountered in RS scenarios in Sec. 6. After an overview of commonly used datasets is provided in Sec. 7, we analyze and compare the state-of-the-art methods in Sec. 8. Finally, we conclude our work and discuss the future directions of oriented object detection in Sec. 9.
@@ 2 From Horizontal Object Detection to Oriented Object Detection
The early object detection methods rely on handcrafted descriptors [95, 27, 32, 165] and machine learning algorithms [16, 3, 82, 2]. These methods often show limited performance due to the weak feature representations. Although lag far behind in accuracy, their instructive insights still have a profound impact on modern detectors, e.g., sliding windows [149, 150], hard negative mining, and bounding box regression [34, 33]. Readers interested in early object detection methods are referred to the recent survey [7] that provides an in-depth analysis of classical object detection methods in RS.
The world has witnessed impressive progress in computer vision with the advance of deep neural networks (DNN) since 2012 [60, 73, 14, 65, 70, 71]. Owing to the persistent enhancement of computing resources, DNN can learn high-level patterns from large-scale datasets in an end-to-end manner. The pioneering studies bring a little glimmer to the object detection field, especially in light of the fact that the performance of handcrafted features-based detectors reached a plateau after 2010. Since then, a growing number of DNN-based detectors have emerged and have dominated the state-of-the-art, thanks to their powerful feature representation.
Early research in the deep learning era is primarily concerned with designing horizontal object detectors [35, 37, 130, 131, 72, 78, 79, 127, 129, 50, 23, 230, 188] for natural scene images taken from a horizontal perspective. Naturally, as horizontal object detectors evolve rapidly, numerous studies are harnessing their immense potential in RS scenarios, e.g., RCNN series [35, 37, 130, 131], YOLO series [127, 129], and RetinaNet [78, 79]. Subsequently, a growing number of efforts pivoted towards refining network structures and crafting innovative data augmentation techniques, all aimed at tackling the core challenges in RS object detection, including scale variation [108, 189, 90, 66], complex background [83, 54, 109], and weak feature responses [145, 168].
Nevertheless, RS images are typically captured from the BEV, leading to objects appearing in arbitrary orientations. Hence, directly applying horizontal object detectors in RS images may encounter the following problems: (1) The intersection-over-union (IoU) between an HBB and the adjacent HBBs can be very large in dense arrangement scenarios, especially for objects with extremely large aspect ratios, as illustrated in Fig. 2b. Thus, the non-maximum suppression (NMS) technique tends to cause missed detection. (2) HBBs are inclined to contain background, whereas OBBs can tightly enclose the objects, achieving more precise localization, as shown in Fig. 2a. Given the above predicament of HBB, OBB is considered more appropriate for RS object detection.
@@
With the remarkable development of detection frameworks [130, 131, 78, 79, 127, 13], backbone networks [65, 85], and robust feature representation [94, 22], the field of object detection has achieved dramatic breakthroughs. Naturally, an intuitive strategy for designing oriented object detectors is to modify representative horizontal object detectors by predicting additional parameters to represent OBB [228]. However, such a straightforward strategy is plagued by several additional challenges, mainly including feature misalignment, spatial misalignment, and OBB regression problems.
@@
(1) Feature Misalignment. The prevailing generic object detectors typically contain a feature extraction network followed by a detection head, where the latter leverages the feature representations generated from the former to make decisions. However, the feature representations are generally extracted via axis-aligned convolutions, thereby exposing non-negligible misalignment with oriented objects, as shown in Fig. 3a. Such a misaligned feature representation degrades the performance of oriented object detectors, due to the lack of rotational information, making the detector struggle to identify objects and regress precise OBBs.
@@
(2) Spatial Misalignment. In addition to feature misalignment, the widely used anchor-based detection methods also struggle with spatial misalignment. Generic anchor-based detectors typically use horizontal anchors as priors thereby having limited overlaps to oriented objects, especially for objects with extremely large aspect ratios, as shown in Fig. 4a. This poses a significant challenge to generic label assignment strategies [130, 131], which assign positive or negative samples depending on the overlaps. Thus, the naïve anchor generation mechanism is likely unable to provide sufficient positive samples during the training process.
@@
(3) OBB Regression Problems. Current detectors commonly utilize the regression paradigm to represent the locations of objects, which has been proven to be effective and yields dramatic achievement. The most commonly used representation methods for OBB include -based and quadrilateral representation. However, the former suffers from the periodicity of angle (PoA), causing the angular boundary discontinuity [200, 203, 125, 126]. Concretely, a small angle difference may cause a large loss change when the angular value approaches the angular boundary range. On the other hand, the later faces challenges related to vertex ordering, because the inappropriate vertex sorting may cause inconsistencies between the vertex sequences of the predicted OBB and the ground truth (GT). Overall, both PoA and vertex ordering problem can seriously confuse the network, leading to training instability. For more details please refer to Sec. A and Sec. B of the Appendix.
@@
To cope with the above dilemmas, various works have been made and achieved notable advancements. Several methods construct well-designed detection frameworks by devising rotated proposal generation networks (Sec. 3.1) or refined heads (Sec. 3.2) to remedy feature misalignment. To deal with spatial misalignment, researchers focus on improving the assignment schemes (Sec. 3.1) or adopting anchor-free mechanisms (Sec. 3.3). As for OBB regression problems, several efforts intends to design effective OBB regression through developing new loss functions (Sec. 4.1) or OBB representation (Sec. 4.2). In addition, high-quality feature representations are crucial for object detection, hence a great deal of effort is concentrated on network designs for better feature representations, including rotation-invariant feature representations (Sec. 5.1) and advanced feature representations (Sec. 5.2). We also cover the solutions for several common issues (Sec. 6) in RS scenarios. Fig. 5 shows the taxonomy of oriented object detection methods in this survey.